7.Calculate Total bill amount for 4 days using the below rates for A flat with 2 adults and 3 children.
1 adult =$ 10 per day
1 Child=$ 5 per day
A. $ 140
B. =$ 60
C. =$ 120
D. =$80

Answer by ikleyn(52813)

.
7.Calculate Total bill amount for 4 days using the below rates for A flat with 2 adults and 3 children.
1 adult =$ 10 per day
1 Child=$ 5 per day
A. $ 140
B. =$ 60
C. =$ 120
D. =$80
~~~~~~~~~~~~~~~~~~~~

Total bill amount = 4 days *(2*10 + 3*5) = 4*(2*10 + 3*5) = 4*35 = 140 dollars.